# Catalog of outbursts of neutron star LMXBs

### v1

@@ Appendix A Outburst details
Swift J1756.9-2508 2009/8 2009/8 5.7e-10 - - 4.4e36 8 Patruno et al. (2009)
Swift J1756.9-2508 2018/4 2018/4 2.5e-10 4.3e22 - 1.9e36 8 Rai & Paul (2019)
Swift J1756.9-2508 2019/6 2019/6 2.6e-10 4.2e22 - 2.0e36 8 Li et al. (2021)
IGR J17591-2342 2018/8 2018/10 3.2e-10 4.4e22 2.4 2.5e36 8 Gusinskaia et al. (2020)
2S 1803-245 1976/5 1976/5 2.2e-8 1.47e22 - 1.4e38 $<$7.3 Cornelisse et al. (2007)
2S 1803-245 1998/4 1998/4 9.1e-9 - - 5.8e37 $<$7.3 Marshall et al. (1998)
@@
SAX J1806.5-2215 2011/2 2017/7 1.5e-9 4.2e22 1.7 1.1e37 $<$8 Del Santo et al. (2012); Shaw et al. (2017), Swift/XRT
MAXI J1807+132 2017/3 2017/3 2.2e-10 - - 4e36 $<$12.4 Shidatsu et al. (2017)
MAXI J1807+132 2019/9 2019/9 1.9e-10 1.3e21 2 3.5e36 $<$12.4 Albayati et al. (2021)
MAXI J1807+132 2023/7 2023/10 2.7e-10 1.3e21 2 4.9e36 $<$12.4 Illiano et al. (2023a), Swift/XRT
SAX J1808.4-3658 1996/9 1996/10 2.2e-9 - - 3.2e36 3.5 in ’t Zand et al. (1998c); Revnivtsev (2003)
SAX J1808.4-3658 1998/4 1998/4 1.3e-9 - - 1.9e36 3.5 Chakrabarty et al. (1998); Bult & van der Klis (2015)
SAX J1808.4-3658 2000/1 2000/4 2.8e-10 1.3e21 2.16 4.1e35 3.5 Wijnands et al. (2001a)
SAX J1808.4-3658 2002/10 2002/11 1.8e-9 - - 2.7e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2005/5 2005/7 1.3e-9 - - 1.9e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2008/9 2008/11 1.2e-9 - - 1.7e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2011/11 2011/11 1.7e-9 - - 2.5e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2015/4 2015/4 1.2e-9 1e21 1.6 1.7e36 3.5 Sanna et al. (2015)
SAX J1808.4-3658 2019/8 2019/8 2.1e-10 2.1e21 2 3.1e35 3.5 Bult et al. (2019b)
SAX J1808.4-3658 2022/8 2022/10 3.7e-10 - - 5.4e35 3.5 Illiano et al. (2023b)
@@
XTE J1810-189 2008/3 2008/6 1.1e-9 3.8e22 1.7 4.0e36 5.5 Chelovekov & Grebenev (2009); Weng et al. (2015)
@@
SAX J1810.8-2609 2007/8 2007/10 8e-10 5e21 2 3e36 $<$5.7 Degenaar et al. (2007); Degenaar & Wijnands (2013a)
@@
XTE J1814-338 2003/6 2003/7 3.7e-10 1.6e21 2.1 2.8e36 8 Papitto et al. (2007); Krauss et al. (2005)
MAXI J1816-195 2022/6 2022/7 3.2e-9 2.4e22 2.6 1.5e37 $<$6.3 Kennea et al. (2022a); Bult et al. (2022b)
Swift J181723.1-164300 2017/8 2017/8 3.6e-10 8e22 1.5 2.1e36 7 Parikh et al. (2017a)
IGR J18245-2452 2013/3 2013/4 1.2e-9 4e21 1.26 4.3e36 5.5 Papitto et al. (2013); Linares et al. (2014)
@@
SAX J1828.5-1037 2008/10 2008/11 1.5e-11 4.1e22 1.6 6.9e34 $<$6.2 Campana (2009); Degenaar et al. (2008), Swift/XRT
SAX J1828.5-1037 2011/11 2011/11 2.9e-10 4.1e22 1.6 1.3e36 $<$6.2 Serino et al. (2016), Swift/XRT
Swift J185003.2-005627 2011/5 2011/7 1.2e-10 1.1e22 1.6 2.0e35 $<$3.7 Degenaar et al. (2012a)
Swift J1858.6-0814 2018/10 2020/5 2.1e-9 1.8e23 - 4.1e37 12.8 Negoro et al. (2020); Hare et al. (2020)
HETE J1900.1-2455 2005/6 2015/10 8e-10 2e21 - 2.1e36 4.7 Degenaar et al. (2017a); Kaaret et al. (2006)
1H 1905+000 1974/11 1985/9 1.5e-10 1.9e21 2.1 1.0e36 7.5 Jonker et al. (2006)
@@
Swift J1922.7-1716 2011/7 2011/8 2.1e-10 2.7e21 2.0 5.8e35 4.8 Degenaar et al. (2012a)
@@
MAXI J1957+032 2016/9 2016/9 6.6e-10 - - 2.0e36 5 Negoro et al. (2016c); Mata Sánchez et al. (2017)
MAXI J1957+032 2022/6 2022/6 8e-10 - - 2.4e36 5 Negoro et al. (2022); Sanna et al. (2022)
XTE J2123-058 1999/6 1999/8 1.7e-9 6e20 2 1.5e37 8.5 Tomsick et al. (1999, 2004)
@@
Note. — Published outbursts of verified neutron star LMXBs. Max $F_{X}$ is highest reported flux in each outburst, transformed to 2-10 keV, generally using the power-law spectral index & $N_{H}$ noted. We assume 1 Crab is Fx(2-10)=2e-8 erg/s/cm2. Peak observed intrinsic $L_{X}$ in 2-10 keV, for the assumed distance (listed). A * at the end date indicates that we do not have evidence that the outburst ended–the last recorded observation found it active.
@@
MXB 1659-298 or MXB 1658-298; the 10 kpc distance estimate is from a PRE burst (Sharma et al., 2018).
@@
XTE J1701-407; the distance of 6.2±$\pm 0.5$ kpc was estimated from a PRE burst (Falanga et al., 2009; Linares et al., 2009). It is not clear whether XTE J1701-407 has ever reached full quiescence since 2016.
@@
4U 1730-22; Bult et al. (2022a) measure the distance of 6.9±$\pm 0.2$ kpc from PRE bursts.
@@
Swift J1734.5-3027 or IGR J17344-3023; Bozzo et al. (2015) measured a distance of 7.2±$\pm 1.5$ kpc from a PRE burst.
@@
KS 1741-293; Chelovekov & Grebenev (2011) place an upper limit of $<$6.2 kpc from analyzing bursts. While this source has generally been seen not to exceed a few $10^{36}$ ergs/s (see e.g. Degenaar & Wijnands 2013b), we notice that the 2011-2012 outburst was reported by Chenevez et al. (2012) using INTEGRAL/JEM-X to reach a 3-10 keV flux of 415±$\pm 75$ mCrab, an order of magnitude brighter. We also identify a Swift/XRT observation on May 14, 2017 that reached 5.2 cts/s, or $L_{X}\sim 6.4*10^{36}$ erg/s.
1A 1744-361, or A 1744-361; its proximity to the Galactic center has led many workers to assume an 8 kpc distance, though the only burst constraint is $<$11 kpc (Galloway et al., 2008). The RXTE/PCA Bulge scans show three likely small outbursts, at MJD 52020, 53281, and 54188, all reaching around $F_{X}=5\times 10^{-11}$ ergs/cm2/s. Given the crowded environment, it would be useful to verify these signals. 1A 1744-361 is not clearly resolved by MAXI, so our sensitivity to its outbursts declines after 2011.
IGR J17445-2747; the proximity of this source to the Galactic Centre suggests an 8 kpc distance. Mereminskiy et al. (2017) place limits of $>$5 and $<$12.3 kpc, from analysis of a burst.
@@
SAX J1747.0-2853: we use the distance of 7.5±$\pm 1.3$ kpc from PRE bursts of Werner et al. (2004). We find a much smaller flux ($6\times 10^{-9}$ ergs/cm2/s, 2-10 keV unabsorbed) in Feb. 2011 than Kennea et al. (2011). The MAXI page notes that the apparent flare in MAXI data in early 2014 is from a different source. It is unclear whether SAX J1747.0-2853 has been active since May 2023.
@@
NGC 6440 X-2 showed numerous weak, short X-ray outbursts in 2009 and 2010, often with a 1-month recurrence time (Heinke et al., 2010) [^10] Note that Heinke et al. (2010) quote $N_{H}=0.59\times 10^{21}$ in their Table 2, but actually used $N_{H}=0.59\times 10^{22}$. The RXTE PCA bulge scan reveals at least 7 faint outbursts during 2009-2011, but does not show other similar features except one in 2007, and possibly 4 in 1999. It is not certain whether NGC 6440 X-2 only produced outbursts in 2007 and 2009-2011, or whether we missed others (due perhaps to changes in the outburst peak flux). We list NGC 6440 X-2’s outbursts in a separate table (Table 2), to quote their MJD dates.
@@
NGC 6440 X-2 2009/7 55040 1.7e-10 2 1.5e36 Heinke et al. (2010), RXTE/PCA
NGC 6440 X-2 2009/8-9 55072 1.6e-10 1.79 1.4e36 Heinke et al. (2010); Altamirano et al. (2010d), RXTE/PCA
@@
NGC 6440 X-2 2010/6 55359 2e-10 - 1.7e36 Patruno & D’Angelo (2013), RXTE/PCA
@@
IGR J17498-2921; Linares et al. (2011) estimate a 7.6 kpc distance from a PRE burst. NICER observations of the 2023 outburst show a rapid drop to quiescence in May 2023.
@@
SAX J1806.5-2215 showed a long outburst, which started in Feb. 2011 (Altamirano et al., 2011a). This was reported in Atels until 2015 (Sguera et al., 2015), but Swift/XRT observations indicate it continued until July 2017, and was no longer visible to Swift/XRT by January 2018. The $<$8 kpc upper limit comes from an analysis of bursts by (Cornelisse et al., 2002a).
@@
SAX J1808.4-3658; Galloway & Cumming (2006) estimated a distance of 3.5$\pm$0.1 kpc, using PRE bursts and the expected mass transfer rate from gravitational radiation.
@@
XTE J1812-182: also known as XMMU J181227.8-181234. This may be the same as 1H1812-182, detected by HEAO-1 with $F_{X}$(2-10 keV)$=7\times 10^{-10}$ erg/cm2/s in 1977/78 (Wood et al., 1984), but the errorbox is large. Goodwin et al. (2019) estimated the distance as 14$\pm$2 kpc, from analysis of a train of several (non-PRE) bursts.
@@ Appendix B Frequently outbursting sources
[…]
@@
4U 1608-52 1970/04 2.0e-8 2.4e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1970/09 1.7e-8 2.1e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1971/06 1.5e-8 1.8e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1971/09 1.7e-8 2.1e37 Lochner & Roussel-Dupre (1994)
@@
4U 1608-52 1977/07 2.4e-8 2.9e37 Clark & Li (1977)
4U 1608-52 1979/01 6.5e-9 7.9e36 Kaluzienski & Holt (1979a)
4U 1608-52 1979/04 1.3e-8 1.6e37 Oda et al. (1979)
@@
4U 1608-52 1983/04 1.6e-8 2.0e37 Mitsuda et al. (1984)
4U 1608-52 1984/07 1.0e-9 1.2e36 Gottwald et al. (1987)
4U 1608-52 1986/03 5.8e-10 7.1e35 Penninx et al. (1989)
@@
4U 1608-52 1996/3 7.1e-9 8.7e36 Méndez et al. (1998)
4U 1608-52 1998/2-4 1.2e-8 1.5e37 Rutledge (2001)
4U 1608-52 2000/01 4.4e-9 5.4e36 Šimon (2004)
@@
4U 1608-52 2002/07 1.7e-8 2.1e37 Šimon (2004); Keek et al. (2008)
4U 1608-52 2003/03 6.3e-9 7.7e36 Miller et al. (2003a)
4U 1608-52 2004/03 5.4e-9 6.6e36 Remillard & Miller (2004)
@@
4U 1608-52 2005/03 2e-8 2.4e37 Keek et al. (2008)
4U 1608-52 2005/09 6.1e-9 7.4e36 Asai et al. (2012)
4U 1608-52 2006/07 1.9e-9 2.3e36 Markwardt et al. (2006); Greco et al. (2006)
4U 1608-52 2007/06-2009/10 7.8e-8 9.5e37 Palmer et al. (2007); Russell et al. (2009)
4U 1608-52 2010/02 6.3e-9 7.7e36 Armas Padilla et al. (2017)
4U 1608-52 2011/03 1.7e-8 2.1e37 Suguzaki et al. (2011)

### v2

@@ Appendix A Outburst details
Swift J1756.9-2508 2009/8 2009/8 5.7e-10 - - 4.4e36 8 Markwardt et al. (2009); Patruno et al. (2009)
Swift J1756.9-2508 2018/4 2018/4 2.5e-10 4.3e22 - 1.9e36 8 Mereminskiy et al. (2018); Rai & Paul (2019)
Swift J1756.9-2508 2019/6 2019/6 2.6e-10 4.2e22 - 2.0e36 8 Sanna et al. (2019); Li et al. (2021)
IGR J17591-2342 2018/8 2018/10 3.2e-10 4.4e22 2.4 2.5e36 8 Ducci et al. (2018); Gusinskaia et al. (2020)
IGR J17597-2201 2001/2 2008/3 4.6e-10 4.5e22 1.7 3.5e36 8 Lutovinov et al. (2003); Markwardt & Swank (2003a); Bodaghee et al. (2012)
2S 1803-245 1976/5 1976/5 2.2e-8 1.47e22 - 1.4e38 $<$7.3 Jernigan (1976); Cornelisse et al. (2007)
2S 1803-245 1998/4 1998/4 9.1e-9 - - 5.8e37 $<$7.3 Muller et al. (1998); Marshall et al. (1998a)
@@
SAX J1806.5-2215 2011/2 2017/7 1.5e-9 4.2e22 1.7 1.1e37 $<$8 Altamirano et al. (2011b); Del Santo et al. (2012); Shaw et al. (2017), Swift/XRT
MAXI J1807+132 2017/3 2017/3 2.2e-10 - - 4e36 $<$12.4 Negoro et al. (2017c); Shidatsu et al. (2017)
MAXI J1807+132 2019/9 2019/9 1.9e-10 1.3e21 2 3.5e36 $<$12.4 Shidatsu et al. (2019); Albayati et al. (2021)
MAXI J1807+132 2023/7 2023/10 2.7e-10 1.3e21 2 4.9e36 $<$12.4 Saikia et al. (2023); Illiano et al. (2023a); Rout et al. (2025), Swift/XRT
SAX J1808.4-3658 1996/9 1996/10 2.2e-9 - - 1.9e36 2.7 in ’t Zand et al. (1998d); Revnivtsev (2003)
SAX J1808.4-3658 1998/4 1998/4 1.3e-9 - - 1.1e36 2.7 Marshall et al. (1998b); Chakrabarty et al. (1998); Bult & van der Klis (2015)
SAX J1808.4-3658 2000/1 2000/4 2.8e-10 1.3e21 2.16 2.4e35 2.7 van der Klis et al. (2000); Wijnands et al. (2001a)
SAX J1808.4-3658 2002/10 2002/11 1.8e-9 - - 1.6e36 2.7 Markwardt et al. (2002a); Bult & van der Klis (2015)
SAX J1808.4-3658 2005/5 2005/7 1.3e-9 - - 1.1e36 2.7 Markwardt et al. (2005); Bult & van der Klis (2015)
SAX J1808.4-3658 2008/9 2008/11 1.2e-9 - - 1.0e36 2.7 Markwardt & Swank (2008b); Bult & van der Klis (2015)
SAX J1808.4-3658 2011/11 2011/11 1.7e-9 - - 1.5e36 2.7 Markwardt et al. (2011); Bult & van der Klis (2015)
SAX J1808.4-3658 2015/4 2015/4 1.2e-9 1e21 1.6 1.0e36 2.7 Sanna et al. (2015a, b)
SAX J1808.4-3658 2019/8 2019/8 2.1e-10 2.1e21 2 1.8e35 2.7 Goodwin et al. (2019b); Russell et al. (2019a); Bult et al. (2019b)
SAX J1808.4-3658 2022/8 2022/10 3.7e-10 - - 3.2e35 2.7 Sanna et al. (2022a); Imai et al. (2022); Illiano et al. (2023b)
@@
XTE J1810-189 2008/3 2008/6 1.1e-9 3.8e22 1.7 4.0e36 5.5 Markwardt & Swank (2008c); Chelovekov & Grebenev (2009); Weng et al. (2015)
@@
SAX J1810.8-2609 2007/8 2007/10 8e-10 5e21 2 3e36 $<$5.7 Parsons et al. (2007); Degenaar et al. (2007); Degenaar & Wijnands (2013a)
@@
XTE J1814-338 2003/6 2003/7 3.7e-10 1.6e21 2.1 2.8e36 8 Markwardt & Swank (2003b); Papitto et al. (2007); Krauss et al. (2005)
MAXI J1816-195 2022/6 2022/7 3.2e-9 2.4e22 2.6 1.5e37 $<$6.3 Negoro et al. (2022a); Kennea et al. (2022a); Bult et al. (2022b)
Swift J181723.1-164300 2017/8 2017/8 3.6e-10 8e22 1.5 2.1e36 7 Barthelmy et al. (2017); Parikh et al. (2017a)
IGR J18245-2452 2013/3 2013/4 1.2e-9 4e21 1.26 4.3e36 5.5 Eckert et al. (2013); Papitto et al. (2013); Linares et al. (2014)
@@
SAX J1828.5-1037 2008/10 2008/11 1.5e-11 4.1e22 1.6 6.9e34 $<$6.2 Campana (2009); Degenaar & Wijnands (2008), Swift/XRT
SAX J1828.5-1037 2011/11 2011/11 2.9e-10 4.1e22 1.6 1.3e36 $<$6.2 Asada et al. (2011); Serino et al. (2016a), Swift/XRT
Swift J185003.2-005627 2011/5 2011/7 1.2e-10 1.1e22 1.6 2.0e35 $<$3.7 Beardmore et al. (2011); Degenaar et al. (2012a)
Swift J1858.6-0814 2018/10 2020/5 2.1e-9 1.8e23 - 4.1e37 12.8 Krimm et al. (2018); Negoro et al. (2020b); Hare et al. (2020)
HETE J1900.1-2455 2005/6 2015/10 8e-10 2e21 - 2.1e36 4.7 Vanderspek et al. (2005); Degenaar et al. (2017a); Kaaret et al. (2006)
1H 1905+000 1974/11 1985/9 1.5e-10 1.9e21 2.1 1.0e36 7.5 Seward et al. (1976); Jonker et al. (2006a)
@@
Swift J1922.7-1716 2011/7 2011/8 2.1e-10 2.7e21 2.0 5.8e35 4.8 Nakahira et al. (2011); Degenaar et al. (2012a)
Swift J1922.7-1716 2024/2 2024/5 4.2e-10 - 2.2 1.2e36 4.8 Nakajima et al. (2024),Swift/XRT
@@
MAXI J1957+032 2016/9 2016/9 6.6e-10 - - 2.0e36 5 Negoro et al. (2016d); Mata Sánchez et al. (2017)
MAXI J1957+032 2022/6 2022/6 8e-10 - - 2.4e36 5 Negoro et al. (2022b); Sanna et al. (2022b)
MAXI J1957+032 2025/5 2025/5 9e-10 - - 3e36 5 Sun et al. (2025); Negoro et al. (2025); Illiano et al. (2025)
XTE J2123-058 1998/6 1998/8 1.7e-9 6e20 2 1.5e37 8.5 Levine et al. (1998); Tomsick et al. (1999, 2004)
@@
Note. — Published outbursts of verified neutron star LMXBs. Max $F_{X}$ is highest reported flux in each outburst, transformed to 2-10 keV, generally using the power-law spectral index & $N_{H}$ noted. We assume 1 Crab is $F_{X}$(2-10)=$2\times 10^{-8}$ erg cm-2 s-1. Peak observed intrinsic $L_{X}$ in 2-10 keV, for the assumed distance (listed). A * at the end date indicates that we do not have evidence that the outburst ended–the last recorded observation found it active. Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
@@
MXB 1659-298 or MXB 1658-298, or MXB 1659-29, or H 1658-298; the 10 kpc distance estimate is from a PRE burst (Sharma et al., 2018).
@@
XTE J1701-407; the distance of 6.2±$\pm 0.5$ kpc was estimated from a PRE burst (Falanga et al., 2009; Linares et al., 2009a). It is not clear whether XTE J1701-407 has ever reached full quiescence since 2016.
@@
4U 1730-22; Bult et al. (2022a) measure the distance of 6.9±$\pm 0.2$ kpc from PRE bursts. The end of its second outburst is uncertain, but it was clearly in quiescence in a Swift/XRT observation in May 2025.
@@
Swift J1734.5-3027 or IGR J17344-3023; Bozzo et al. (2015a) measured a distance of 7.2±$\pm 1.5$ kpc from a PRE burst.
@@
KS 1741-293; Chelovekov & Grebenev (2011) place an upper limit of $<$6.2 kpc from analyzing bursts. While this source has generally been seen not to exceed a few $10^{36}$ ergs s-1 (see e.g. Degenaar & Wijnands 2013b), we notice that the 2011-2012 outburst was reported by Chenevez et al. (2012) using INTEGRAL/JEM-X to reach a 3-10 keV flux of 415±$\pm 75$ mCrab, an order of magnitude brighter. We also identify a Swift/XRT observation on May 14, 2017 that reached 5.2 cts/s, or $L_{X}\sim 6.4*10^{36}$ erg/s.
1A 1744-361, or A 1744-361; its proximity to the Galactic center has led many workers to assume an 8 kpc distance, though the only burst constraint is $<$11 kpc (Galloway et al., 2008). The RXTE/PCA Bulge scans show three likely small outbursts, at MJD 52020, 53281, and 54188, all reaching around $F_{X}=5\times 10^{-11}$ ergs cm-2 s-1. Given the crowded environment, it would be useful to verify these signals. 1A 1744-361 is not clearly resolved by MAXI, so our sensitivity to its outbursts declines after 2011.
IGR J17445-2747; the proximity of this source to the Galactic Centre suggests an 8 kpc distance. Mereminskiy et al. (2017b) place limits of $>$5 and $<$12.3 kpc, from analysis of a burst.
@@
SAX J1747.0-2853: we use the distance of 7.5±$\pm 1.3$ kpc from PRE bursts of Werner et al. (2004). We find a much smaller flux ($6\times 10^{-9}$ ergs cm-2 s-1, 2-10 keV unabsorbed) in Feb. 2011 than Kennea et al. (2011). The MAXI page notes that the apparent flare in MAXI data in early 2014 is from a different source.
@@
NGC 6440 X-2 showed numerous weak, short X-ray outbursts in 2009 and 2010, often with a 1-month recurrence time (Heinke et al., 2010) [^11] Note that Heinke et al. (2010) quote $N_{H}=0.59\times 10^{21}$ cm-2 in their Table 2, but actually used $N_{H}=0.59\times 10^{22}$ cm-2.. The RXTE PCA bulge scan reveals at least 7 faint outbursts during 2009-2011, but does not show other similar features except one in 2007, and possibly 4 in 1999. It is not certain whether NGC 6440 X-2 only produced outbursts in 2007 and 2009-2011, or whether we missed others (due perhaps to changes in the outburst peak flux). We list NGC 6440 X-2’s outbursts in a separate table (Table 2), to quote their MJD dates.
@@
NGC 6440 X-2 2009/7 55040 1.7e-10 2 1.5e36 Heinke & Budac (2009); Heinke et al. (2010), RXTE/PCA
NGC 6440 X-2 2009/8-9 55072 1.6e-10 1.79 1.4e36 Heinke et al. (2009, 2010); Altamirano et al. (2010d), RXTE/PCA
@@
NGC 6440 X-2 2010/6 55359 2e-10 - 1.7e36 Patruno et al. (2010); Patruno & D’Angelo (2013), RXTE/PCA
@@
IGR J17498-2921; Linares et al. (2011b) estimate a 7.6 kpc distance from a PRE burst, but Galloway et al. (2024) infer $5.7^{+0.6}_{-0.5}$ kpc from more detailed modeling, which we adopt. NICER observations of the 2023 outburst show a rapid drop to quiescence in May 2023.
@@
IGR J17597-2201, or XTE J1759-220; the 8 kpc distance estimate assumes it is near the Galactic Center, which is consistent with the (unconstraining, $<16$ kpc) X-ray burst constraint (Galloway et al., 2008).
@@
SAX J1806.5-2215 showed a long outburst, which started in Feb. 2011 (Altamirano et al., 2011b). This was reported in Atels until 2015 (Sguera et al., 2015), but Swift/XRT observations indicate it continued until July 2017, and was no longer visible to Swift/XRT by January 2018. The $<$8 kpc upper limit comes from an analysis of bursts by (Cornelisse et al., 2002a).
@@
SAX J1808.4-3658; Galloway & Cumming (2006) estimated a distance of 3.5$\pm$0.1 kpc, using PRE bursts; this was revised to $2.7\pm 0.3$ kpc by Galloway et al. (2024).
@@
XTE J1812-182: also known as XMMU J181227.8-181234. This may be the same as 1H1812-182, detected by HEAO-1 with $F_{X}$(2-10 keV)$=7\times 10^{-10}$ erg/cm2/s in 1977/78 (Wood et al., 1984), but the errorbox is large. Goodwin et al. (2019a) estimated the distance as 14$\pm$2 kpc, from analysis of a train of several (non-PRE) bursts.
@@ Appendix B Frequently outbursting sources
4U 1608-52, or 4U 1608-522, X1608-52, QX Nor, in Table 3. We assume a distance of 3.2$\pm$0.3 kpc (Galloway et al., 2008). This is reasonably consistent with the estimate of Özel et al. (2016) of $3.6\pm 0.3$ kpc. 4U 1608-52 has remained continually in low-level outburst for several years at several points in its history, most notably 1999-2000 (Wachter et al., 2002) and 2007-2009 (Šimon, 2020), but also probably in the 1970s (Šimon, 2020). We see 4U 1608-52 to be in a continuous outburst from July 2022 through July 2023, with a normal, bright, short outburst superimposed in November 2022.
@@
4U 1608-52 1970/4 2.0e-8 2.4e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1970/9 1.7e-8 2.1e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1971/6 1.5e-8 1.8e37 Matilsky et al. (1972); Lochner & Roussel-Dupre (1994)
4U 1608-52 1971/9 1.7e-8 2.1e37 Lochner & Roussel-Dupre (1994); Tananbaum et al. (1976)
@@
4U 1608-52 1977/7 2.4e-8 2.9e37 Clark & Li (1977)
4U 1608-52 1979/1 6.5e-9 7.9e36 Kaluzienski & Holt (1979a)
4U 1608-52 1979/4 1.3e-8 1.6e37 Oda et al. (1979)
@@
4U 1608-52 1983/4 1.6e-8 2.0e37 Mitsuda et al. (1984)
4U 1608-52 1984/7 1.0e-9 1.2e36 Gottwald et al. (1987)
4U 1608-52 1986/3 5.8e-10 7.1e35 Penninx et al. (1989)
@@
4U 1608-52 1996/3 7.1e-9 8.7e36 Marshall & Angelini (1996); Méndez et al. (1998)
4U 1608-52 1998/2-4 1.2e-8 1.5e37 Cui & Swank (1998); Rutledge (2001)
4U 1608-52 2000/1 4.4e-9 5.4e36 Šimon (2004)
@@
4U 1608-52 2002/7 1.7e-8 2.1e37 Šimon (2004); Keek et al. (2008)
4U 1608-52 2003/3 6.3e-9 7.7e36 Miller et al. (2003a)
4U 1608-52 2004/3 5.4e-9 6.6e36 Remillard & Miller (2004)
@@
4U 1608-52 2005/3 2e-8 2.4e37 Keek et al. (2008)
4U 1608-52 2005/9 6.1e-9 7.4e36 Asai et al. (2012)
4U 1608-52 2006/7 1.9e-9 2.3e36 Markwardt et al. (2006); Greco et al. (2006)
4U 1608-52 2007/6-2009/10 7.8e-8 9.5e37 Palmer et al. (2007); Russell et al. (2009)
4U 1608-52 2010/2 6.3e-9 7.7e36 Morii (2010); Armas Padilla et al. (2017)
4U 1608-52 2011/3 1.7e-8 2.1e37 Suguzaki et al. (2011); Asai et al. (2013)